OpenStudy (anonymous):

Will give a medal and fan(if wanted)~ A sandbag was thrown downward from a building. The function f(t) = -16t2 - 64t + 512 shows the height f(t), in feet, of the sandbag after t seconds. Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? Part C: Use your answer in part B to determine the axis of symmetry for f(x)? ( I think I got the answers, just need help to make sure they're correct)

OpenStudy (anonymous):

I got this for Part A Part A: f(t) = -16(x^2 + 4x - 32) = -16(x + 8)(x - 4) The x-intercept of the function is x = -8 and x = 4

OpenStudy (anonymous):

And this for Part B Part B: f(t) = -16(x^2 + 4x + 4) + 64 + 256 = -16(x + 2)^2 + 320 The vertex is (-2,320)

OpenStudy (anonymous):

And for Part C Part C: The axis of symmetry is x = -2

OpenStudy (anonymous):

@undeadknight26 May you help me? n.n

undeadknight26 (undeadknight26):

Sounds bout right ._.

OpenStudy (anonymous):

xD I hate asking questions so I am sorry for this :P

OpenStudy (anonymous):

Just wanna make sure it's correct; hopefully it is o ~o

undeadknight26 (undeadknight26):

Its alright little chibi <3333 Your ratio is still beautiful...Keep being kawaii <3

OpenStudy (anonymous):

Lol :P Well, I guess I'll find out sooner or later~ It's okay though..my grade is good o_o.. Thanks for the help xD